OpenStudy (anonymous):

Given a triangle with b=4, c=10, and A=84*, what is the length of a? Round to the nearest tenth. I know you have to use the law of cosines, but i am really confused on how to do this.

OpenStudy (kc_kennylau):

\[a^2=b^2+c^2-2bc\cos(A)\] \[a^2=116-80\cos(84)\] \[a=10.375\]